Find the characteristic equation of the matrix A =

2 1 1

0 1 0

1 1 2

 and hence find the matrix

represented by A8 − 5A7 + 7A6 − 3A5 + A4 − 5A3 + 8A2 − 2A + I.


1
Expert's answer
2021-07-07T09:01:45-0400

Characteristic equation "\det(A-\lambda I)=0"


"\begin{vmatrix}
 2-\lambda & 1 & 1 \\
 0 & 1-\lambda & 0 \\
 1 & 1 & 2-\lambda \\
\end{vmatrix}=0"


"(2-\lambda)\begin{vmatrix}
 1-\lambda & 0 \\
 1 & 2-\lambda
\end{vmatrix}-\begin{vmatrix}
 0 & 0 \\
 1 & 2-\lambda
\end{vmatrix}+\begin{vmatrix}
 0 & 1-\lambda \\
 1 & 1
\end{vmatrix}=0"

"(2-\lambda)^2(1-\lambda)-(1-\lambda)=0"

"(1-\lambda)(4-4\lambda+\lambda^2-1)=0"

"(1-\lambda)^2(3-\lambda)=0"

"\lambda_1=1, \lambda_2=1, \lambda_3=3"

The equation can be written as


"\lambda^3-5\lambda^2+7\lambda-3=0"

According to Cayley Hamilton theorem, every matrix is the root of it's eigen matrix. Then


"A^3-5A^2+7A-3=0"

Given sum

This sum can be written as,


"A^8 \u2212 5A^7 + 7A^6 \u2212 3A^5"




"+ A^4 \u2212 5A^3 + 8A^2 \u2212 2A + I"

"= (A^3 - 5A^2+7A -3)(A^5+A) + (A^2+A+I)"

Since "A^3-5A^2+7A-3=0," then


"A^8 \u2212 5A^7 + 7A^6 \u2212 3A^5"




"+ A^4 \u2212 5A^3 + 8A^2 \u2212 2A + I"

"=A^2+A+I"

"A^2=\begin{bmatrix}
 2 & 1 & 1 \\
 0 & 1 & 0 \\
 1 & 1 & 2 \\
\end{bmatrix}\begin{bmatrix}
 2 & 1 & 1 \\
 0 & 1 & 0 \\
 1 & 1 & 2 \\
\end{bmatrix}"

"=\begin{bmatrix}
 4+0+1 & 2+1+1 & 2+0+2 \\
 0+0+0 & 0+1+0 & 0+0+0 \\
 2+0+2 & 1+1+2 & 1+0+4 \\
\end{bmatrix}"

"=\begin{bmatrix}
 5 & 4 & 4 \\
 0 & 1 & 0 \\
 4 & 4 & 5 \\
\end{bmatrix}"


"A^8 \u2212 5A^7 + 7A^6 \u2212 3A^5"




"+ A^4 \u2212 5A^3 + 8A^2 \u2212 2A + I"

"=A^2+A+I"

"=\begin{bmatrix}
 5 & 4 & 4 \\
 0 & 1 & 0 \\
 4 & 4 & 5 \\
\end{bmatrix}+\begin{bmatrix}
 2 & 1 & 1 \\
 0 & 1 & 0 \\
 1 & 1 & 2 \\
\end{bmatrix}+\begin{bmatrix}
 1 & 0 & 0 \\
 0 & 1 & 0 \\
 0 & 0 & 1 \\
\end{bmatrix}"


"=\begin{bmatrix}
 8 & 5 & 5 \\
 0 & 3 & 0 \\
 5 & 5 & 8 \\
\end{bmatrix}"